Assuming the mean and SD are the same for this year’s national exam, what percentage of test-takers scored above and below you if your score is 1900 and given that X ~ N(1600, 140). Using the empirical rule, find the percentage of students who scored between 1880 and 1920?


1
Expert's answer
2021-12-15T17:33:37-0500

X ~ N(1600, 140)

Population mean = 1600

Variance = 140

"\sigma=\sqrt{variance}=\sqrt{140}=11.83"

"P(X<1900)=P(Z<\frac{1900-1600}{11.83})"

"P(X<1900)=P(Z<25.36)"

Using normal distribution table, we find that probability is:

"P(X<1900)=1≈100\%"

This means 100 percent of test-takers scored above and below if score is 1900.


As per the empirical rule, the percentage of students who scored between 1880 and 1920 is:

The empirical rule formula said that;

68% of data falls within 1 standard deviation from the mean - that means between "\mu-\sigma"  and "\mu+\sigma"

95% of data falls within 2 standard deviations from the mean - between "\mu-2\sigma"  and "\mu+2\sigma"

99.7% of data falls within 3 standard deviations from the mean - between "\mu-3\sigma" and "\mu+3\sigma"

Given that;

Population mean = 1600

Variance = 140

"\sigma=\sqrt{variance}=\sqrt{140}=11.83"

"(\u03bc\u2212\u03c3,\u03bc+\u03c3)\u2248(1588.2,1611.8)"

"(\u03bc\u22122\u03c3,\u03bc+2\u03c3)\u2248(1576.3,1623.7)"

"(\u03bc\u22123\u03c3,\u03bc+3\u03c3)\u2248(1564.5,1635.5)"